# Navigator: A Decentralized Scheduler for Latency-Sensitive ML Workflows

### 27 Feb 2024 v1

@@ 5.1.2 Data Transfer between Task Dispatchers
Although Navigator’s algorithms anticipate object transfers and include cost estimates, Navigator itself does not currently initiate such actions. Instead, the MLs themselves access objects by performing application-layer actions such as opening a file that is in fact hosted by Cascade, reading a key-value object hosted by Cascade, or receiving a pub-sub object that actually is a Cascade object. In all of these cases, Cascade itself will then transparently fetch the object either from a local host cache, from local storage if the server making the access is a home node for the object, or from one of the home nodes if the object is hosted on other servers. Navigator does not currently try to "peek" into the Cascade host cache, but instead adopts the view that every object accessed during an ML job will be in memory "somewhere" in the system, and also that if the object in question is a large ML model - the case that can become costly - that DMA transfer from the local cache into GPU memory is comparable in speed to RDMA or DPDK transfer from a remote server’s memory to the local GPU memory.
Navigator treats one kind of object differently: results of a task that become inputs to a successor task. Here, the distinction is that whereas the outputs of some programs need to be saved and will be reused later, when we consider a DFG there may also be transient outputs produced by one stage and then consumed and discarded by its successor stages. Here, the cost depends on where the producer stage runs, and where the consumers run. If an object is small, or if it is large but resides in the GPU cache on a node where the consumer will run, this cost is 0. This form of co-location is clearly very beneficial. Otherwise, Navigator simply assumes that a DMA or RDMA transfer will occur, and models the cost accordingly. The resulting model is presented in Figure 4.
@@ 5.2 Shared State Table Considerations
Figure 5: Navigator Shared State Table (SST)
While most objects stored in Cascade are simply key-value tuples with home nodes, but accessible from anywhere, the SST is a very different kind of storage structure. As summarized above, the SST is an extensible table replicated over all servers. Navigator’s "instance" of the SST has one row per server, and a fixed row format. The resulting replicated object is an inherently $O(n^{2})$ data structure: every node has a copy of the data structure, and when a node updates its row and pushes that update, one RDMA write per peer will be issued, to update the corresponding row in that peer. Fortunately, $n$ in our target settings would generally be fairly small: most edge compute clusters have just a few (2 to perhaps 64) blade-style computers in a rack. Even so, we limit the update rate to ensure that overhead is not excessive.
@@
Navigator’s decentralized algorithm requires that a node making a scheduling choice be able to estimate the states of other nodes, but does not require perfect accuracy or lock-step coordination. Accordingly, we decided to limit the frequency of SST updates using a parameter in the Navigator configuration file, and then to experimentally determine the lowest acceptable update rate, thereby ensuring that SST overheads will be minimized. In Section 6.2 we report on these experiments, confirming that Navigator performs well even with somewhat stale metadata, and justifying an update frequency of 5 pushes per second: negligible given that RDMA can send 75M small messages per second on each network link.
@@ 5.3 GPU Memory Management
Navigator’s GPU Memory Manager controls caching for recently used ML models in GPU memory. It performs model fetching and eviction as new tasks are assigned to the associated worker. Its fetching and eviction policy can be configured by users, depending on the application. We implemented and experimented with two management policies: FIFO and queue-lookahead.
@@ 5.3.2 Queue-Lookahead
Upon eviction, GPU Memory Manager looks ahead into the execution queue, examining some fixed number of future tasks and their required models, and giving higher priority to models that will be used sooner. Then it sorts the list of models in GPU memory based on the priority. If a model must be evicted to make room for the model of the current task, GPU Memory Manager evicts models from the lowest priority to the highest. This way Navigator can avoid evicting models that are needed for the near-future tasks, and the associated overhead of fetching these models soon again.
@@ 5.4 Simulation
Although Navigator is a fully working system, we wanted a way to explore scenarios beyond what we can reasonably deploy on our limited-scale testbed. Accordingly, we implemented an event-driven simulator [git repo]. The simulator itself is similar to that of Sparrow [13]. It models all elements of the environment described in Section 3, and supports the same request types as the real system, as illustrated in Figure 1. When estimating network delays and execution times, the simulator uses values we measured in the real system for input/output sizes, the sizes of ML models for our workload, and the task duration. Navigator’s simulator models the task arrival, task waiting on the queue, task execution, and task dispatch as events, processed in the order of the simulated time. We validated the simulator against the real experiment using the same scale of 5 workers and the same workload, and observed that the performance difference between simulation and real experiment lies within 5% of the median numeric values.
@@ 6 Experiments and Evaluation
We evaluate Navigator on a dedicated cluster, using 5 servers as the worker nodes and one as a client node. All workers are Dell PowerEdge R740 machines with dual Intel Xeon Gold 6242 processors and 192 GB memory. Each is additionally equipped with a Tesla T4 GPU having 16GB memory. The servers are connected by InfiniBand with RDMA-enabled 100Gbps network. Experiments in Section 6.2 and 6.3.2 are evaluated on a real system. Experiment in Section 6.5 is performed using the simulator.
@@ 6.2 Scheduling Algorithm Analysis
Two primary considerations arise when scheduling workflows with DFG dependencies. One entails optimizing for dependencies within each job instance. For example, tasks that could be processed in parallel can be sent to different workers to maximize the level of concurrency during execution; tasks with large intermediate data dependencies can be grouped to run on a single worker to avoid data transfer delays. The other entails optimization reflecting the current state of the system as a whole, as represented in the shared metadata table. Here, decisions will avoid overloading a worker that is already at its peak capacity. Both approaches offer potential benefits, but they lead to very different schedules. The Navigator algorithm takes both factors into consideration, using a hybrid scheduling scheme. Below in Section 6.2.1 we describe two scheduling schemes that each focus on one factor. Then Section 6.2.2 compares Navigator with these two baseline options.
@@ 6.2.2 Performance Comparison Among Schedulers
Figure 5(a) shows the low-load case. Clients send pipeline processing requests at the average rate of 0.5 requests per second, with a Poison distribution on request types. The box plot shows the distribution of slow_down_factors for different instances, broken down by job category. The top and bottom of the box represent the first and third quartile of the data, respectively. The whiskers represent 1.5 times the upper and lower quartiles, and the dots show outliers. While all schedulers perform well under this case with close to optimal slow_down_factor, Navigator is the closest to the optimal schedule (slow_down_factor of 1.0).
Figure 5(b) repeats the experiment with a job arrival rate having a mean of 2 requests per second under a Poison distribution. This places the cluster under more pressure, so the slow_down_factor for all four scheduling schemes increases due to the higher load. Navigator continues to outperform, yielding overall performance 2x to 4x faster than HEFT and Hash for the translation and Q&A pipelines, and 20x to 30x faster for the image description and 3D perception pipeline. The more extreme slow_down_factors in image description and 3D perception pipelines are because of their relatively short runtimes compared to translation and Q&A, which made these pipelines more susceptible to the system overheads caused by sub-optimal schedules.
Figure 5(c) experiments with different request rates under a Poison distribution. The plot shows the average slow_down_factors for a mixed workload. Navigator has the closest to ideal performance at different request rates. The HEFT algorithm fails to adjust to the higher system loads due to not taking the workers’ load into consideration. JIT performs better than the HEFT and Hash schemes due to its in-time task assignment mechanism, but not as well as Navigator because of it lacks a mechanism for intra-job coordination.
Table 1 shows average latency and other GPU-related performance metrics for the experiment presented in Figure 5(b). GPU utilization is the percentage of time during which the GPU was actively processing computations. Memory utilization is the percentage of GPU memory being used, an indicator of memory-pressure on the GPU. Energy consumption is the total amount of energy consumed when running the full workloads. The GPU cache hit rate is the percentage of times when the required model is already in the GPU memory. The results indicate that Navigator consumes similar GPU resources and energy to other schedulers, yet achieves superior end-to-end latency.
@@ 6.3 Ablation Analysis
Navigator employs multiple features that contribute to scheduling. To understand the importance of each, we conduct an ablation analysis by selectively disabling features.
@@ 6.3.1 Importance of Scheduler Features
To show the effect of using dynamic adjustment, this experiment disables dynamic adjustment that Navigator normally employs as each job instance executes, only retaining the preliminary planning phase. Figure 7 reveals a significant performance degradation.
@@
The scheduler prioritizes workers whose GPU cache already contains the required models, as reflected in the metadata table. As shown in Figure 7, Navigator incurs an 8x degradation without this mechanism. The GPU cache hit rate also decreases to 90% (as opposed to 99% when this optimization is enabled).
@@ 6.3.2 Sensitivity Analysis
Recall from Section 3 that we made a decision to limit the update rate for the SST-hosted metadata on which Navigator depends, forcing the system to run with potentially stale data. The question arises of how sensitive Navigator’s scheduling quality actually is as a function of the degree of staleness.
@@
The graph tells us quite a bit more, however. Because ML model fetching to the GPU memory is an infrequent event, we see that Navigator can tolerate a much higher level of staleness for GPU cache-content information than for load. The delay for the latter has a significant impact beyond the threshold of 200ms (corresponding to 5 pushes per second). To make this concrete, at the 200ms rate in a cluster of 32 nodes, each participant would send and receive 160 RDMA updates per second (out of an RDMA rate limit of 75M/second): a trivially low overhead.
@@ 6.4 Production trace
Our next experiment considers scheduling performance for the workflows described in Figure 1 using a public trace from the Alibaba production GPU cluster [4, 63, 85]. We rescale this trace to adapt it to the capacity of our experimental system. Figure 8(a) shows request arrival rates on a timeline, while Figures 8(b)-8(e) show completion times as a function of arrival times. The plots make it clear that the Hash scheduling scheme is least tolerant of bursts of job instances, at least in a small worker cluster. Navigator has the best completion times even with these unpredictable request patterns.
@@ 6.5 Scalability Analysis
Hashing is widely used for scalability in today’s large clusters [43, 75, 26], under the assumption that it leads to a more even distribution of workloads and full use of resources. Our next experiment shows that in fact, Navigator uses hardware resources more efficiently. For this work we create a large-scale experiment using a simulation study. The workload is Poisson with a mean of 40 requests per second. We consider a range of up to 250 workers and examine a series of cluster utilization metrics as well as the median slow_down_factor over the job instances of all four types in Figure 1.
@@
As seen in Figure 10, Hashing is effective in the sense that it utilizes all available workers in each round, and the median slow_down_factor decreases as the number of workers increases, due to its inherent load-balancing property, approaching its lower bound at the scale of around 100 workers. Navigator prioritizes the workers with models in GPU memory and expands the worker set only if by doing so, mean job completion times would be improved. As observed in the experiment with the same workload, the slow_down_factor reaches its lower bound with just 50 active workers: half the resources required for the Hashing scheduler! These unused machines could be put into power-saving mode, reducing overall platform power consumption. With very large numbers of workers, Hashing slightly outperforms Navigator (at 150 machines and beyond), but to gain this tiny benefit it keeps three times as many machines active.
@@ 7 Related work
Scheduling for workflows with inter-job dependencies has been explored in many prior efforts  [48, 16, 8, 49]. Wing [48] focuses on data-analytic queries; Apollo [16] focuses on data-processing jobs with dependencies in production cluster; HEFT [8] schedules for general-purpose computation workflow on heterogeneous processors. Batchy [49] is a general-purpose dataflow graph scheduler that uses batch-processing profiles to balance the efficiency and latency under SLO constraints. The estimation-based job planning and deferred correction mechanism used in Apollo [16] resembles the scheduling scheme in Navigator; while the task assignment algorithm in HEFT [8] resembles Navigator’s job planning algorithm. However, direct application of these methods to the ML workflows in Navigator does not give latency-optimal scheduling result because they do not consider ML model re-use, GPU memory consumption and collocation of ML models.
Navigator’s graphical workflows are similar to those considered in  [21, 38, 18, 53, 74]. Video storm[21] and Nexus [38] schedule pipeline-structure workflows for video stream. Orion [74] schedules serverless DAGs with right-sizing, bundling and pre-warming techniques to meet latency SLOs. Stream Processing Engines (SPEs) scheduler [18] is similar with Orion [74] in that both use workload prediction model to estimate the worker waittime and provision the workers accordingly. These provisioning methods rely on workload prediction, but potentially sacrifice latency when the workload distribution drifts or if a burst of jobs puts the system under resource stress. InferLine[53] introduces a low-frequency combinatorial planner and a high-frequency auto-scaling tuner mechanism that it uses to upscale and downscale the number of workers as load varies. The process of worker selection for provision and up-scaling is less addressed in InferLine [53], which becomes nontrivial when the model sizes are significantly larger and when there are only limited GPU memories. Navigator addresses this scenario by scheduler and resource management co-design, which optimizes the task assignment while balancing the co-location of ML model objects in Navigator-managed GPU memory.
@@ 8 Conclusion
Navigator manages memory and schedules tasks in support of latency-sensitive ML workflows structured as DFGs. A novel scheduling-aware eviction policy yields high GPU cache hit rates, avoiding unnecessary model fetching and eviction. Our evaluation compared Navigator with state of the art schedulers under a variety of request rates and workload patterns, showing 2x to 6x speedup with similar resource use, while leaving excess servers completely idle. Navigator also turns out to perform well with respect to energy consumption.

### 28 Feb 2024 v2

@@ 5.1.2 Data Transfer between Task Dispatchers
Although Compass’s algorithms anticipate object transfers and include cost estimates, Compass itself does not currently initiate such actions. Instead, the MLs themselves access objects by performing application-layer actions such as opening a file that is in fact hosted by Cascade, reading a key-value object hosted by Cascade, or receiving a pub-sub object that actually is a Cascade object. In all of these cases, Cascade itself will then transparently fetch the object either from a local host cache, from local storage if the server making the access is a home node for the object, or from one of the home nodes if the object is hosted on other servers. Compass does not currently try to "peek" into the Cascade host cache, but instead adopts the view that every object accessed during an ML job will be in memory "somewhere" in the system, and also that if the object in question is a large ML model - the case that can become costly - that DMA transfer from the local cache into GPU memory is comparable in speed to RDMA or DPDK transfer from a remote server’s memory to the local GPU memory.
Compass treats one kind of object differently: results of a task that become inputs to a successor task. Here, the distinction is that whereas the outputs of some programs need to be saved and will be reused later, when we consider a DFG there may also be transient outputs produced by one stage and then consumed and discarded by its successor stages. Here, the cost depends on where the producer stage runs, and where the consumers run. If an object is small, or if it is large but resides in the GPU cache on a node where the consumer will run, this cost is 0. This form of co-location is clearly very beneficial. Otherwise, Compass simply assumes that a DMA or RDMA transfer will occur, and models the cost accordingly. The resulting model is presented in Figure 4.
@@ 5.2 Shared State Table Considerations
Figure 5: Compass Shared State Table (SST)
While most objects stored in Cascade are simply key-value tuples with home nodes, but accessible from anywhere, the SST is a very different kind of storage structure. As summarized above, the SST is an extensible table replicated over all servers. Compass’s "instance" of the SST has one row per server, and a fixed row format. The resulting replicated object is an inherently $O(n^{2})$ data structure: every node has a copy of the data structure, and when a node updates its row and pushes that update, one RDMA write per peer will be issued, to update the corresponding row in that peer. Fortunately, $n$ in our target settings would generally be fairly small: most edge compute clusters have just a few (2 to perhaps 64) blade-style computers in a rack. Even so, we limit the update rate to ensure that overhead is not excessive.
@@
Compass’s decentralized algorithm requires that a node making a scheduling choice be able to estimate the states of other nodes, but does not require perfect accuracy or lock-step coordination. Accordingly, we decided to limit the frequency of SST updates using a parameter in the Compass configuration file, and then to experimentally determine the lowest acceptable update rate, thereby ensuring that SST overheads will be minimized. In Section 6.2 we report on these experiments, confirming that Compass performs well even with somewhat stale metadata, and justifying an update frequency of 5 pushes per second: negligible given that RDMA can send 75M small messages per second on each network link.
@@ 5.3 GPU Memory Management
Compass’s GPU Memory Manager controls caching for recently used ML models in GPU memory. It performs model fetching and eviction as new tasks are assigned to the associated worker. Its fetching and eviction policy can be configured by users, depending on the application. We implemented and experimented with two management policies: FIFO and queue-lookahead.
@@ 5.3.2 Queue-Lookahead
Upon eviction, GPU Memory Manager looks ahead into the execution queue, examining some fixed number of future tasks and their required models, and giving higher priority to models that will be used sooner. Then it sorts the list of models in GPU memory based on the priority. If a model must be evicted to make room for the model of the current task, GPU Memory Manager evicts models from the lowest priority to the highest. This way Compass can avoid evicting models that are needed for the near-future tasks, and the associated overhead of fetching these models soon again.
@@ 5.4 Simulation
Although Compass is a fully working system, we wanted a way to explore scenarios beyond what we can reasonably deploy on our limited-scale testbed. Accordingly, we implemented an event-driven simulator [git repo]. The simulator itself is similar to that of Sparrow [13]. It models all elements of the environment described in Section 3, and supports the same request types as the real system, as illustrated in Figure 1. When estimating network delays and execution times, the simulator uses values we measured in the real system for input/output sizes, the sizes of ML models for our workload, and the task duration. Compass’s simulator models the task arrival, task waiting on the queue, task execution, and task dispatch as events, processed in the order of the simulated time. We validated the simulator against the real experiment using the same scale of 5 workers and the same workload, and observed that the performance difference between simulation and real experiment lies within 5% of the median numeric values.
@@ 6 Experiments and Evaluation
We evaluate Compass on a dedicated cluster, using 5 servers as the worker nodes and one as a client node. All workers are Dell PowerEdge R740 machines with dual Intel Xeon Gold 6242 processors and 192 GB memory. Each is additionally equipped with a Tesla T4 GPU having 16GB memory. The servers are connected by InfiniBand with RDMA-enabled 100Gbps network. Experiments in Section 6.2 and 6.3.2 are evaluated on a real system. Experiment in Section 6.5 is performed using the simulator.
@@ 6.2 Scheduling Algorithm Analysis
Two primary considerations arise when scheduling workflows with DFG dependencies. One entails optimizing for dependencies within each job instance. For example, tasks that could be processed in parallel can be sent to different workers to maximize the level of concurrency during execution; tasks with large intermediate data dependencies can be grouped to run on a single worker to avoid data transfer delays. The other entails optimization reflecting the current state of the system as a whole, as represented in the shared metadata table. Here, decisions will avoid overloading a worker that is already at its peak capacity. Both approaches offer potential benefits, but they lead to very different schedules. The Compass algorithm takes both factors into consideration, using a hybrid scheduling scheme. Below in Section 6.2.1 we describe two scheduling schemes that each focus on one factor. Then Section 6.2.2 compares Compass with these two baseline options.
@@ 6.2.2 Performance Comparison Among Schedulers
Figure 5(a) shows the low-load case. Clients send pipeline processing requests at the average rate of 0.5 requests per second, with a Poison distribution on request types. The box plot shows the distribution of slow_down_factors for different instances, broken down by job category. The top and bottom of the box represent the first and third quartile of the data, respectively. The whiskers represent 1.5 times the upper and lower quartiles, and the dots show outliers. While all schedulers perform well under this case with close to optimal slow_down_factor, Compass is the closest to the optimal schedule (slow_down_factor of 1.0).
Figure 5(b) repeats the experiment with a job arrival rate having a mean of 2 requests per second under a Poison distribution. This places the cluster under more pressure, so the slow_down_factor for all four scheduling schemes increases due to the higher load. Compass continues to outperform, yielding overall performance 2x to 4x faster than HEFT and Hash for the translation and Q&A pipelines, and 20x to 30x faster for the image description and 3D perception pipeline. The more extreme slow_down_factors in image description and 3D perception pipelines are because of their relatively short runtimes compared to translation and Q&A, which made these pipelines more susceptible to the system overheads caused by sub-optimal schedules.
Figure 5(c) experiments with different request rates under a Poison distribution. The plot shows the average slow_down_factors for a mixed workload. Compass has the closest to ideal performance at different request rates. The HEFT algorithm fails to adjust to the higher system loads due to not taking the workers’ load into consideration. JIT performs better than the HEFT and Hash schemes due to its in-time task assignment mechanism, but not as well as Compass because of it lacks a mechanism for intra-job coordination.
Table 1 shows average latency and other GPU-related performance metrics for the experiment presented in Figure 5(b). GPU utilization is the percentage of time during which the GPU was actively processing computations. Memory utilization is the percentage of GPU memory being used, an indicator of memory-pressure on the GPU. Energy consumption is the total amount of energy consumed when running the full workloads. The GPU cache hit rate is the percentage of times when the required model is already in the GPU memory. The results indicate that Compass consumes similar GPU resources and energy to other schedulers, yet achieves superior end-to-end latency.
@@ 6.3 Ablation Analysis
Compass employs multiple features that contribute to scheduling. To understand the importance of each, we conduct an ablation analysis by selectively disabling features.
@@ 6.3.1 Importance of Scheduler Features
To show the effect of using dynamic adjustment, this experiment disables dynamic adjustment that Compass normally employs as each job instance executes, only retaining the preliminary planning phase. Figure 7 reveals a significant performance degradation.
@@
The scheduler prioritizes workers whose GPU cache already contains the required models, as reflected in the metadata table. As shown in Figure 7, Compass incurs an 8x degradation without this mechanism. The GPU cache hit rate also decreases to 90% (as opposed to 99% when this optimization is enabled).
@@ 6.3.2 Sensitivity Analysis
Recall from Section 3 that we made a decision to limit the update rate for the SST-hosted metadata on which Compass depends, forcing the system to run with potentially stale data. The question arises of how sensitive Compass’s scheduling quality actually is as a function of the degree of staleness.
@@
The graph tells us quite a bit more, however. Because ML model fetching to the GPU memory is an infrequent event, we see that Compass can tolerate a much higher level of staleness for GPU cache-content information than for load. The delay for the latter has a significant impact beyond the threshold of 200ms (corresponding to 5 pushes per second). To make this concrete, at the 200ms rate in a cluster of 32 nodes, each participant would send and receive 160 RDMA updates per second (out of an RDMA rate limit of 75M/second): a trivially low overhead.
@@ 6.4 Production trace
Our next experiment considers scheduling performance for the workflows described in Figure 1 using a public trace from the Alibaba production GPU cluster [4, 63, 85]. We rescale this trace to adapt it to the capacity of our experimental system. Figure 8(a) shows request arrival rates on a timeline, while Figures 8(b)-8(e) show completion times as a function of arrival times. The plots make it clear that the Hash scheduling scheme is least tolerant of bursts of job instances, at least in a small worker cluster. Compass has the best completion times even with these unpredictable request patterns.
@@ 6.5 Scalability Analysis
Hashing is widely used for scalability in today’s large clusters [43, 75, 26], under the assumption that it leads to a more even distribution of workloads and full use of resources. Our next experiment shows that in fact, Compass uses hardware resources more efficiently. For this work we create a large-scale experiment using a simulation study. The workload is Poisson with a mean of 40 requests per second. We consider a range of up to 250 workers and examine a series of cluster utilization metrics as well as the median slow_down_factor over the job instances of all four types in Figure 1.
@@
As seen in Figure 10, Hashing is effective in the sense that it utilizes all available workers in each round, and the median slow_down_factor decreases as the number of workers increases, due to its inherent load-balancing property, approaching its lower bound at the scale of around 100 workers. Compass prioritizes the workers with models in GPU memory and expands the worker set only if by doing so, mean job completion times would be improved. As observed in the experiment with the same workload, the slow_down_factor reaches its lower bound with just 50 active workers: half the resources required for the Hashing scheduler! These unused machines could be put into power-saving mode, reducing overall platform power consumption. With very large numbers of workers, Hashing slightly outperforms Compass (at 150 machines and beyond), but to gain this tiny benefit it keeps three times as many machines active.
@@ 7 Related work
Scheduling for workflows with inter-job dependencies has been explored in many prior efforts  [48, 16, 8, 49]. Wing [48] focuses on data-analytic queries; Apollo [16] focuses on data-processing jobs with dependencies in production cluster; HEFT [8] schedules for general-purpose computation workflow on heterogeneous processors. Batchy [49] is a general-purpose dataflow graph scheduler that uses batch-processing profiles to balance the efficiency and latency under SLO constraints. The estimation-based job planning and deferred correction mechanism used in Apollo [16] resembles the scheduling scheme in Compass; while the task assignment algorithm in HEFT [8] resembles Compass’s job planning algorithm. However, direct application of these methods to the ML workflows in Compass does not give latency-optimal scheduling result because they do not consider ML model re-use, GPU memory consumption and collocation of ML models.
Compass’s graphical workflows are similar to those considered in  [21, 38, 18, 53, 74]. Video storm[21] and Nexus [38] schedule pipeline-structure workflows for video stream. Orion [74] schedules serverless DAGs with right-sizing, bundling and pre-warming techniques to meet latency SLOs. Stream Processing Engines (SPEs) scheduler [18] is similar with Orion [74] in that both use workload prediction model to estimate the worker waittime and provision the workers accordingly. These provisioning methods rely on workload prediction, but potentially sacrifice latency when the workload distribution drifts or if a burst of jobs puts the system under resource stress. InferLine[53] introduces a low-frequency combinatorial planner and a high-frequency auto-scaling tuner mechanism that it uses to upscale and downscale the number of workers as load varies. The process of worker selection for provision and up-scaling is less addressed in InferLine [53], which becomes nontrivial when the model sizes are significantly larger and when there are only limited GPU memories. Compass addresses this scenario by scheduler and resource management co-design, which optimizes the task assignment while balancing the co-location of ML model objects in Compass-managed GPU memory.
@@ 8 Conclusion
Compass manages memory and schedules tasks in support of latency-sensitive ML workflows structured as DFGs. A novel scheduling-aware eviction policy yields high GPU cache hit rates, avoiding unnecessary model fetching and eviction. Our evaluation compared Compass with state of the art schedulers under a variety of request rates and workload patterns, showing 2x to 6x speedup with similar resource use, while leaving excess servers completely idle. Compass also turns out to perform well with respect to energy consumption.